The sound level measured in a room by a person watching a movie on a home theater system varies from 50 dB

 during a quiet part to 90 dB

 during a loud part. Approximately how many times louder is the latter sound?


1
Expert's answer
2021-09-21T18:21:00-0400

The sound intensity level "\beta" of a sound wave is defined by the equation


"\beta = (10 \,dB) \log{\cfrac{I}{I_0}}"


Then, since we have two sounds ("\beta_1=50\,dB;\beta_2=90\,dB") we substitute to find the intensity of the loud part I2:


"\beta_2-\beta_1 = (10 \,dB) \Bigg( \log{\cfrac{I_2}{I_0}}-\log{\cfrac{I_1}{I_0}} \Bigg)
\\ \beta_2-\beta_1 = (10 \,dB) \Big( \log{{I_2}}-\log{{I_0}}-\log{{I_1}}+\log{{I_0}} \Big)
\\ \beta_2-\beta_1 = (10 \,dB) \big( \log{{I_2}}-\log{{I_1}} \big)"



"\cfrac{{I_2}}{{I_1}} = \Large{10}^{\frac{{\beta_2-\beta_1 }}{{(10 \,dB)}}}=\Large{10}^{\frac{{(90-50)dB }}{{(10 \,dB)}}}=\Large{10}^{4}"


"\Large \implies I_2=10^4\cdot I_1"


In conclusion, I2 (the intensity of the loud part) is approximately 104 times bigger than I1 (the intensity of the quiet part).


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.